An individual deposits $100 per month into an account which pays interest at a rate of 6% per year compounded monthly. The value of the account after five years is closest to:

a. $5,630

b. $6,210

c. $6,764

d. $6,977
Answer #1

Answer

Option d

The formula is:

A=monthly payment =100

r=monthly interest rate =6/12=0.5%

n=number of months=12*5=60

FV=future value in the account=?

then,

FV=100*\frac{(1+0.005)^{60}-1}{0.005}=
6977.00305

The amount in the deposit is $6977.00305.